# Corresponding Author

### v3

@@ 6.2 Personalized Video Generation
With the rising popularity of video generation xing2024survey , video personalization has also begun to attract attention. In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
Appearance-based video personalization. This task focuses on transferring subject appearance from static images to video sequences. The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis. The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning 145wang2024customvideo , data augmentation 138chen2023videodreamer ; 145wang2024customvideo , and attention manipulation 83jiang2023videobooth ; 212liu2024stylecrafter ; 145wang2024customvideo . Additionally, several studies 82zhao2023videoassembler ; 83jiang2023videobooth ; 212liu2024stylecrafter ; 105feng2023dreamoving have explored the PTA framework. These diffusion models are specifically tailored to synthesize videos based on the image references.
Motion-based video personalization. In this task, the reference input switches to a video clip containing a consistent action. A common approach is to fine-tune the video diffusion model by reconstructing the action clip 150wu2023tune ; 103song2023save ; 140jeong2023vmc ; 148esser2023structure ; 136he2023animate ; 102wu2023lamp ; 104materzynska2023customizing . However, distinguishing between appearance and motion within the reference video can be challenging. To solve this problem, SAVE 103song2023save applies appearance learning to ensure that appearance is excluded from the motion learning phase. Additionally, VMC 140jeong2023vmc removes the background information during training prompt construction.
Appearance and motion personalization. When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously. MotionDirector 101zhao2023motiondirector utilizes spatial and temporal losses to facilitate learning across these dimensions. Another approach, DreamVideo 142wei2023dreamvideo , incorporates residual features from a randomly selected frame to emphasize subject information. This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
@@ 6.3 Personalized 3D generation
Personalized 3D Generation refers to the process of creating customized 3D models or scenes based on 2D SoI images. Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods. This tuned model then guides the optimization of a 3D Neural Radiance Field (NeRF) model mildenhall2021nerf for each specific prompt 120lin2023magic3d ; 126shi2023mvdream . In the second phase, DreamFusion poole2022dreamfusion is the key technique, which introduces Score Distillation Sampling (SDS) to train a 3D model capable of rendering images aligned with 2D diffusion model. Building on this foundation, several methods have been developed to improve the workflow. DreamBooth3D 121raj2023dreambooth3d structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement. Consist3D 127ouyang2023chasing enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization. TextureDreamer 133yeh2024texturedreamer focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
Additionally, advancements extend to 3D avatar rendering and dynamic scenes. Animate124 128zhao2023animate124 and Dream-in-4D 130zheng2023unified integrate video diffusion for 4D dynamic scene support within the 3D optimization process. In avatar rendering, PAS 122azadi2023text generates 3D body poses configurable by avatar settings, StyleAvatar3D 123zhang2023styleavatar3d facilitates 3D avatar generation based on images, and AvatarBooth 124zeng2023avatarbooth employs dual fine-tuned diffusion models for separate face and body generation.
@@ 6.4 Attack and Defense
This rapid advancement raises concerns about the ethical implications of PCS, particularly in areas such as misinformation, privacy violations, and the creation of deepfakes. There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception. To mitigate this, Anti-DreamBooth 13van2023anti aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results. The basic idea is to maximize the reconstruction loss of the surrogate model. Additionally, 50zhang2023backdooring suggests to predefine a collection of trigger words and meaningless images. These data are paired and incorporated during the training phase. Once the trigger words are encountered, the synthesized image will be intentionally altered for safeguarding.
@@ 6.5 Other Emerging Directions
Several works are exploring different personalization extensions. For example, SVG personalization is introduced by 36zhang2023text , in which a parameter-efficient fine-tuning method is applied to create SVGs. After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output. Another application, 360-degree panorama customization 25wang2024customizing , is also emerging as a potential tool for personalization in the digital imaging realm.
@@ 7.1 Evaluation Dataset
DreamBench 2ruiz2023dreambooth serves as the primary evaluation benchmark for DreamBooth 2ruiz2023dreambooth , containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
DreamBench-v2 11chen2024subject expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
Custom-10 6kumari2023multi used in Custom Diffusion 6kumari2023multi evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
Custom-101 6kumari2023multi is the latest released dataset by the authors of Custom Diffusion 6kumari2023multi , which comprises 101 subjects to provide a broader scope of evaluation.
Stellar 95achlioptas2023stellar specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
@@ 7.2 Evaluation Metrics
CLIP-T measures semantic alignment using the cosine similarity between CLIP radford2021learning embeddings of generated images and their text prompts.
ImageReward xu2023imagereward , HPS Score (v1/v2) wu2023human ; wu2023humanv2 , and PickScore kirstain2023pick employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
@@
CLIP-I evaluates subject preservation through CLIP image embedding similarity between generations and references. Optimal values balance fidelity (high scores) against overfitting (excessive scores that ignore text guidance).
DINO-I caron2021emerging provides complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
Fréchet Inception Distance (FID) FIDheusel2017gans quantifies the statistical similarity between generated and real image distributions through Inception-V3.
Inception Score (IS) ISsalimans2016improved evaluates both image quality and diversity through classifier confidence.
In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS 76yeh2023navigating introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality. Stellar 95achlioptas2023stellar develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score. These metrics ensure a structured and detailed evaluation of personalized models. This evolving metric landscape reflects the growing sophistication of PCS systems, with optimal evaluation typically requiring combinations of general and task-specific measures.
@@ 7.3 New Benchmark on SoTA Methods
Although multiple evaluation datasets have been proposed in this area, there is still an urgent need for a standardized benchmark that systematically assesses performance across diverse PCS methodologies.To address this limitation, we present a comprehensive evaluation built upon existing works.
For Object. We collect a dataset of 47 subjects from available resources 1gal2022image ; 2ruiz2023dreambooth ; 6kumari2023multi . Following the methodology of DreamBooth 2ruiz2023dreambooth , we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity. Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects. To evaluate the performance, we utilize the text prompts from DreamBooth 2ruiz2023dreambooth . This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
For Face. We collect 15 subjects from Celeb-A liu2018large dataset. We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
Settings. To assess existing representative PCS methods using our constructed test dataset, we compute CLIP-T to assess the text alignment. We calculate CLIP-I for subject fidelity in object generation. For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN zhang2016mtcnn and calculate the pairwise identity similarity using FaceNet schroff2015facenet . To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation. We generate 4 images for each test prompt and set the same random seed for all methods.
Results. The evaluation results are shown in Tab.1. It is evident that no method excels simultaneously in both visual fidelity and text alignment metrics. This highlights a significant challenge currently faced by PCS methods: achieving the optimal trade-off between subject preservation and editability. Striking this balance is difficult, as high subject fidelity often comes at the cost of prompt fidelity and vice versa. Furthermore, we note that higher visual fidelity does not always equate to better performance. The generated images sometimes exhibit patterns that closely mirror the reference images and ignore the prompt guidance. This phenomenon primarily arises from model overfitting on the reference input, which hinders the model’s ability to generalize. Consequently, the visual similarity metric yields higher results based on the mirrored output and the reference, rather than providing an accurate representation of the performance.
Table 1: Evaluation results of representative methods on our evaluation dataset.
@@
Object Textual Inversion 1gal2022image TTF SD 1.5 0.199 0.749
Dreambooth 2ruiz2023dreambooth TTF SD 1.5 0.286 0.772
P+ 23voynov2023p+ TTF SD 1.4 0.244 0.643
Custom Diffusion 6kumari2023multi TTF SD 1.4 0.307 0.722
NeTI 61alaluf2023neural TTF SD 1.4 0.283 0.801
SVDiff 8han2023svdiff TTF SD 1.5 0.282 0.776
Perfusion 7tewel2023key TTF SD 1.5 0.273 0.691
ELITE 19wei2023elite PTA SD 1.4 0.292 0.765
BLIP-Diffusion 10li2024blip PTA SD 1.5 0.292 0.772
IP-Adapter 40ye2023ip PTA SD 1.5 0.2722 0.825
SSR Encoder 221zhang2024ssr PTA SD 1.5 0.288 0.792
MoMA song2024moma PTA SD 1.5 0.322 0.748
Diptych Prompting 209shin2024large PTA FLUX 1.0 dev 0.327 0.722
$\lambda$-eclipse 222patel2024lambda PTA Kandinsky 2.2 0.272 0.824
MS-Diffusion 220personalizationms PTA SDXL 0.298 0.777
Face CrossInitialization 66pang2023cross TTF SD 2.1 0.261 0.469
Face2Diffusion 170shiohara2024face2diffusion PTA SD 1.4 0.265 0.588
SSR Encoder 221zhang2024ssr PTA SD 1.5 0.233 0.490
FastComposer 54xiao2023fastcomposer PTA SD 1.5 0.230 0.516
IP-Adapter 40ye2023ip PTA SD 1.5 0.292 0.462
IP-Adapter 40ye2023ip PTA SDXL 0.292 0.642
PhotoMaker 28li2023photomaker PTA SDXL 0.311 0.547
InstantID 37wang2024instantid PTA SDXL 0.278 0.707
@@ 8.1 Overfitting Problem
As discussed in Sec. 7.3, current PCS systems face a critical challenge of overfitting because of a limited set of reference images. This overfitting problem manifests in two ways: 1) Loss of SoI editability. The personalized model tends to produce images that rigidly mirror the SoI in the reference, such as consistently depicting a cat in an identical pose. 2) Irrelevant semantic inclusion. The irrelevant elements in the references are generated in the output, such as backgrounds or objects not pertinent to the current context.
To investigate the rationale behind, Compositional Inversion 9zhang2023compositional observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words. This is also found in another work 66pang2023cross that the learnable token embeddings deviate significantly from the distribution of the initial embedding. In addition, there is evidence 9zhang2023compositional ; 24jia2023taming ; 20hao2023vico suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
To address this issue, many solutions have been proposed. Most methods discussed in Sec. 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation. However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance 9zhang2023compositional or the context prompt has a similar semantic correlation with the irrelevant elements in the reference 151zhang2024generative . It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments. Therefore, an effective strategy and robust evaluation metrics are urgently needed to achieve broader adoption and greater satisfaction in practical applications.
@@ 8.2 Trade-off on Subject Fidelity and Text Alignment
The ultimate goal of personalized content synthesis is to create systems that not only render the SoI with high fidelity but also effectively respond to textual prompts. However, achieving excellence in both areas simultaneously presents a notable conflict. Specifically, high subject fidelity typically involves capturing and reproducing detailed and specific features of the SoI. This often requires the model to minimize the reconstruction loss to replicate delicate characteristics accurately. Conversely, text alignment necessitates that the system flexibly adapts the SoI according to varying textual descriptions. These descriptions may suggest changes in pose, expression, environment, or stylistic alterations that do not aim to reconstruct the exact visualization in the reference. As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details. To address this inherent conflict, Perfusion 7tewel2023key proposes to regularize the attention projections by these two items. 35cao2023concept decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment. Despite these efforts, there still remains room for further exploration and refinement of this issue. Enhanced model architectures, innovative training methodologies, and more dynamic data handling strategies could potentially provide new pathways to better balance the demands of subject and text fidelity in PCS systems.

### v4

@@ 6.2 Personalized Video Generation
With the rising popularity of video generation [123], video personalization has also begun to attract attention. In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
Appearance-based video personalization. This task focuses on transferring subject appearance from static images to video sequences. The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis. The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning [81], data augmentation [80, 81], and attention manipulation [124, 54, 81]. Additionally, several studies [125, 124, 54, 126] have explored the PTA framework. These diffusion models are specifically tailored to synthesize videos based on the image references.
Motion-based video personalization. In this task, the reference input switches to a video clip containing a consistent action. A common approach is to fine-tune the video diffusion model by reconstructing the action clip [127, 128, 129, 130, 131, 132, 133]. However, distinguishing between appearance and motion within the reference video can be challenging. To solve this problem, SAVE [128] applies appearance learning to ensure that appearance is excluded from the motion learning phase. Additionally, VMC [129] removes the background information during training prompt construction.
Appearance and motion personalization. When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously. MotionDirector [134] utilizes spatial and temporal losses to facilitate learning across these dimensions. Another approach, DreamVideo [135], incorporates residual features from a randomly selected frame to emphasize subject information. This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
@@ 6.3 Personalized 3D generation
Personalized 3D Generation refers to the process of creating customized 3D models or scenes based on 2D SoI images. Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods. This tuned model then utilize Score Distillation Sampling (SDS) [136] to train a 3D Neural Radiance Field (NeRF) model [137] for each specific prompt [138, 139]. Building on this foundation, several methods have been developed to improve the workflow. DreamBooth3D [140] structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement. Consist3D [141] enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization. TextureDreamer [142] focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
Additionally, advancements extend to 3D avatar rendering and dynamic scenes. Animate124 [143] and Dream-in-4D [144] integrate video diffusion for 4D dynamic scene support within the 3D optimization process. In avatar rendering, PAS [145] generates 3D body poses configurable by avatar settings, StyleAvatar3D [146] facilitates 3D avatar generation based on images, and AvatarBooth [147] employs dual fine-tuned diffusion models for separate face and body generation.
@@ 6.4 Attack and Defense
This rapid advancement raises concerns about the ethical implications of PCS, particularly in areas such as misinformation, privacy violations, and the creation of deepfakes. There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception. To mitigate this, Anti-DreamBooth [148] aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results. The basic idea is to maximize the reconstruction loss of the surrogate model. Additionally, [149] suggests to predefine a collection of trigger words and meaningless images. These data are paired and incorporated during the training phase. Once the trigger words are encountered, the synthesized image will be intentionally altered for safeguarding.
@@ 6.5 Other Emerging Directions
Several works are exploring different personalization extensions. For example, SVG personalization is introduced by [150], in which a parameter-efficient fine-tuning method is applied to create SVGs. After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output. Another application, 360-degree panorama customization [151], is also emerging as a potential tool for personalization in the digital imaging realm.
@@ 7.1 Evaluation Dataset
DreamBench [4] serves as the primary evaluation benchmark for DreamBooth [4], containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
DreamBench-v2 [52] expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
Custom-10 [20] used in Custom Diffusion [20] evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
Custom-101 [20] is the latest released dataset by the authors of Custom Diffusion [20], which comprises 101 subjects to provide a broader scope of evaluation.
Stellar [35] specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
@@ 7.2 Evaluation Metrics
CLIP-T measures semantic alignment using the cosine similarity between CLIP [42] embeddings of generated images and their text prompts.
ImageReward [152], HPS Score (v1/v2) [153, 154], and PickScore [155] employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
@@
CLIP-I evaluates subject preservation through CLIP image embedding similarity between generations and references. Optimal values should balance fidelity (high scores) against overfitting (excessive scores that ignore text guidance).
DINO-I [156] provides a complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
Fréchet Inception Distance (FID) [157] quantifies the statistical similarity between generated and real image distributions through Inception-V3 [158].
In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS [36] introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality. Stellar [35] develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score. These metrics ensure a structured and detailed evaluation of personalized models. This evolving metric landscape reflects the growing sophistication of PCS systems, with optimal evaluation typically requiring combinations of general and task-specific measures.
@@ 7.3 New Benchmark on SoTA Methods
Although multiple evaluation datasets have been proposed in this area, there is still an urgent need for a standardized benchmark that systematically assesses performance across diverse PCS methodologies. To address this limitation, we present a new test dataset called Persona with a comprehensive evaluation built upon existing works.
For Object. Persona includes 47 subjects from available resources [7, 4, 20]. Following the methodology of DreamBooth [4], we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity. Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects. To evaluate the performance, we utilize the text prompts from DreamBooth [4]. This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
For Face. We also collected 15 subjects from Celeb-A [159] into our Persona dataset. We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
Settings. To assess existing representative PCS methods using our constructed test dataset, we compute CLIP-T to assess the text alignment. We calculate CLIP-I for subject fidelity in object generation. For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN [160] and calculate the pairwise identity similarity using FaceNet [161]. To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation. We generate 4 images for each test prompt and set the same random seed for all methods.
Results. The evaluation results are shown in Table 1. It is evident that no method excels simultaneously in both visual fidelity and text alignment metrics. This highlights a significant challenge currently faced by PCS methods: achieving the optimal trade-off between subject preservation and editability. Striking this balance is difficult, as high subject fidelity often comes at the cost of prompt fidelity and vice versa. Furthermore, we note that higher visual fidelity does not always equate to better performance. The generated images sometimes exhibit patterns that closely mirror the reference images and ignore the prompt guidance. This phenomenon primarily arises from model overfitting on the reference input, which hinders the model’s ability to generalize. Consequently, the visual similarity metric yields higher results based on the mirrored output and the reference, rather than providing an accurate representation of the performance.
Table 1: Evaluation results of representative methods on our Persona evaluation dataset.
@@
Object Textual Inversion [7] TTF SD 1.5 0.199 0.749
Dreambooth [4] TTF SD 1.5 0.286 0.772
P+ [18] TTF SD 1.4 0.244 0.643
Custom Diffusion [20] TTF SD 1.4 0.307 0.722
NeTI [30] TTF SD 1.4 0.283 0.801
SVDiff [22] TTF SD 1.5 0.282 0.776
Perfusion [21] TTF SD 1.5 0.273 0.691
ELITE [47] PTA SD 1.4 0.292 0.765
BLIP-Diffusion [57] PTA SD 1.5 0.292 0.772
IP-Adapter [51] PTA SD 1.5 0.2722 0.825
SSR Encoder [55] PTA SD 1.5 0.288 0.792
MoMA [59] PTA SD 1.5 0.322 0.748
Diptych Prompting [101] PTA FLUX 1.0 dev 0.327 0.722
$\lambda$-eclipse [112] PTA Kandinsky 2.2 0.272 0.824
MS-Diffusion [56] PTA SDXL 0.298 0.777
Face CrossInitialization [33] TTF SD 2.1 0.261 0.469
Face2Diffusion [162] PTA SD 1.4 0.265 0.588
SSR Encoder [55] PTA SD 1.5 0.233 0.490
FastComposer [5] PTA SD 1.5 0.230 0.516
IP-Adapter [51] PTA SD 1.5 0.292 0.462
IP-Adapter [51] PTA SDXL 0.292 0.642
PhotoMaker [46] PTA SDXL 0.311 0.547
InstantID [6] PTA SDXL 0.278 0.707
@@ 8.1 Overfitting Problem
As discussed in Section 7.3, current PCS systems face a critical challenge of overfitting because of a limited set of reference images. This overfitting problem manifests in two ways: 1) Loss of SoI editability. The personalized model tends to produce images that rigidly mirror the SoI in the reference, such as consistently depicting a cat in an identical pose. 2) Irrelevant semantic inclusion. The irrelevant elements in the references are generated in the output, such as backgrounds or objects not pertinent to the current context.
To investigate the rationale behind, Compositional Inversion [87] observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words. This is also found in another work [33] that the learnable token embeddings deviate significantly from the distribution of the initial embedding. In addition, there is evidence [87, 76, 71] suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
To address this issue, many solutions have been proposed. Most methods discussed in Section 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation. However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance [87] or the context prompt has a similar semantic correlation with the irrelevant elements in the reference [19]. It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments. Therefore, an effective strategy and robust evaluation metrics are urgently needed to achieve broader adoption and greater satisfaction in practical applications.
@@ 8.2 Trade-off on Subject Fidelity and Text Alignment
The ultimate goal of personalized content synthesis is to create systems that not only render the SoI with high fidelity but also effectively respond to textual prompts. However, achieving excellence in both areas simultaneously presents a notable conflict. Specifically, high subject fidelity typically involves capturing and reproducing detailed and specific features of the SoI. This often requires the model to minimize the reconstruction loss to replicate delicate characteristics accurately. Conversely, text alignment necessitates that the system flexibly adapts the SoI according to varying textual descriptions. These descriptions may suggest changes in pose, expression, environment, or stylistic alterations that do not aim to reconstruct the exact visualization in the reference. As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details. To address this inherent conflict, Perfusion [21] proposes to regularize the attention projections by these two items. [163] decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment. Despite these efforts, there still remains room for further exploration and refinement of this issue. Enhanced model architectures, innovative training methodologies, and more dynamic data handling strategies could potentially provide new pathways to better balance the demands of subject and text fidelity in PCS systems.